OpenStudy (anonymous):

Sue counted the turtles and birds in the pond. She counted 14 heads and 44 legs. How many turtles did Sue count?

OpenStudy (anonymous):

4 legs for each turtle (x) 2 legs for each bird (y) 44 legs total x+y=14 4x+2y=44 (-2) (x+y=14) = -2x-2y=-28 -2x-2y=-28 4x+2y=44 ^ add these 2 equations together 2x=16 x=8 Now substitute back into the original problem x+y=14 8+y=14 -8 -8 y=6 8 turtles, 6 birds